Journalist Report – May 15th

Today was Sol 10.

This morning there was an EVA taken by two crew members who went to Marble ritual. The aim of the EVA was to continue on with drone testing. However, during the testing of the search and recover function with a dummy aeroplane, there was a take-off failure.

After the EVA, maintenance work was done on the Hab. The water pump was changed to prevent further leaking. After the installation the pressure increased enough to activate the hot water in the boiler. In addition the water filter was changed.

There was progress made on the processing of the all the photos of the pelican nebula and the robotic telescope finished the observations on the Western Veil Nebula. Poor weather prevented further observations in the solar telescopes. Work progresses on revising the rover, a final attempt will be made with a 2am conference tonight with the UK.

The crew continued on with their group problem solving game.

For outreach a space outfit was made for a Barbie to encourage women in STEM fields with a university in the UK.

Finally, the crew prepared homemade bread and pizza, harvesting tomatoes for both bruschetta and margarita pizza.

Copyright © The Mars Society. All rights reserved. | Main Site